OpenStudy (anonymous):

Find the exact value of the trigonometric function. csc(-630 degrees)

OpenStudy (abhisar):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (abhisar):

Csc(-630°) = -csc(630°)= -csc(2*360°+90°)

OpenStudy (abhisar):

=-csc(90°)=-1

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

thankss :D

OpenStudy (abhisar):

Did you know you can give medals to people who help you with questions? Just click the \(\color{blue}{\text{blue "best response" button}}\) to show appreciation for the people who help you.

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)

OpenStudy (anonymous):

can you answer this too cos 210

OpenStudy (abhisar):

\(\bf Steps\) 1 - If the angle x is negative, we first use a formula for negative angles such as sin (- x) = - sin (x), cos (- x) = cos (x) and so on. 2 - Next we locate the terminal side of the angle in question, directly or using a positive coterminal angle t, which gives the sign of the trigonometric function. 3 - We find the reference angle Tr to the angle in question and use that fact that f(x) = + or - f(Tr). The sign + or - is determined using the quadrant found in step 2. If the angle in question or its coterminal angle are in quadrant 1, this last step is not needed.

OpenStudy (anonymous):

Or you could see your Unit Circle and search where Cosine is 210.

OpenStudy (anonymous):

oh anyway in what quadrant is csc -630 ?

OpenStudy (abhisar):

first quadrant

OpenStudy (abhisar):

its 2*360 + 90

OpenStudy (abhisar):



OpenStudy (anonymous):

it its in the quadrant 1 . why is the answer is negative ?

OpenStudy (anonymous):

if its in the quadrant 1 . why is the answer is negative ?

OpenStudy (abhisar):

because the question has -630°

OpenStudy (abhisar):

csc(-630°)=-csc(630°) by negative angle theorem (point 1)

OpenStudy (abhisar):

csc(630°) will be equal to 1, then we will put -ve sign on it.

OpenStudy (abhisar):

getting it ?

OpenStudy (anonymous):

ah okay. can you give me a tips of solving the exact value because we have a test tommorow about that

OpenStudy (abhisar):

ohkay i'll show u how to find cos 210°

OpenStudy (abhisar):

cos 210° = cos (270°-60°)....agree ?

OpenStudy (abhisar):

ohkay leave this..i'll start a new post and explain how to find values.

OpenStudy (anonymous):

oh i used method is 210-180

OpenStudy (anonymous):

okay i wait